# Construction Identification and Disambiguation Using BERT: A Case Study of NPN

### v1

@@ Abstract
Construction Grammar hypothesizes that knowledge of a language consists chiefly of knowledge of form–meaning pairs (“constructions”) that include vocabulary, general grammar rules, and even idiosyncratic patterns. Recent work has shown that transformer language models represent at least some constructional patterns, including ones where the construction is rare overall. In this work, we probe BERT’s representation of the form and meaning of a minor construction of English, the NPN (noun–preposition–noun) construction—exhibited in such expressions as face to face and day to day—which is known to be polysemous. We construct a benchmark dataset of semantically annotated corpus instances (including distractors that superficially resemble the construction). With this dataset, we train and evaluate probing classifiers. They achieve decent discrimination of the construction from distractors, as well as sense disambiguation among true instances of the construction, revealing that BERT embeddings carry indications of the construction’s semantics. Moreover, artificially permuting the word order of true construction instances causes them to be rejected, indicating sensitivity to matters of form. We conclude that BERT does latently encode at least some knowledge of the NPN construction going beyond a surface syntactic pattern and lexical cues.
@@ 1 Introduction
Thus, there is need to evaluate language models on a range of linguistic tasks which go beyond the more studied “core” linguistic phenomena. Indeed, beyond the more mainstream notions of linguistic structure and information, there is also work on investigating LM knowledge of more idiosyncratic constructions, as defined by Construction Grammar. Construction Grammar is broadly a family of linguistic theories which consider all parts of language to be made up of constructions, which are pairings of linguistic forms with meaning or function (Goldberg 1995; Croft 2001, inter alia). It remains unclear the extent to which LMs may implicitly view constructions as distinct units. A substantial and growing amount of research has recently focused on the intersection of LM knowledge and Construction Grammar (Tayyar Madabushi et al., 2020; Tseng et al., 2022; Pannitto and Herbelot, 2023; Veenboer and Bloem, 2023, inter alia), with a particular focus on argument structure constructions (Li et al., 2022), the English Comparative Correlative (Weissweiler et al., 2022), and the English AANN construction (Chronis et al., 2023; Mahowald, 2023). While these studies have provided valuable insight into LM processing of constructions with varying levels of schematicity, there remain many constructions which have not been addressed at all in previous work. Furthermore, while Zhou et al. (2024) do test model understanding of constructions which are similar in form, no past work has focused on individual constructions as polysemous units. We argue this is a gap in past work, as constructions, like words, can have related but distinct meanings that must be properly disambiguated in context in order for correct interpretation.
@@ 2 The NPN Construction
Given the general rules of English, the NPN construction has several unique properties. Firstly, the nouns almost always lack determiners, which is unusual for count nouns like “day”. Secondly, the construction can occur in a variety of syntactic positions, including as an adverbial modifier (as in section 2) and as a prenominal modifier (as in section 2). Finally, it conveys a meaning which is not entirely predictable from its components, and varies considerably depending on the preposition. Common meanings of the NPN construction are the succession meaning (shown in section 2) and the matching/comparison meaning (shown in section 2). See Jackendoff (2008) for an overview of the NPN construction and the common meanings associated with various prepositional lemmas.
@@
While there are arguably examples of NPNs where the two nouns are not identical, we limit our analysis to cases where the two nouns in the construction match exactly.
@@ 3.1 Corpus Gathering and Cleaning
After extracting all sentences which contained a possible instance of NtoN, we then manually clean the data, removing sentences that were either too short (<5 tokens) or contained too many typos. We annotate all instances of the construction for their semantic subtype, and double annotate roughly 25% of the dataset, achieving an agreement of 84% between the two annotators. 111Disagreements between the two annotators were resolved through discussion and a gold label was chosen jointly. The final dataset has 6599 instances of NtoN, of which 1885 were double annotated.
@@ 3.2 Near Minimal Pairs
These cases are not instances of the NtoN construction, but they do provide a set of negative examples which we can use to probe the model’s ability to recognize true NtoN constructions. Throughout this paper, we refer to this set of examples as instances of the NtoN distractors, since we test of if the model is “distracted” by the shallow similarity of the examples to the NPN construction. We refer to true examples of NtoN as instances of the NtoN construction. Since these NtoN examples exhibit the same surface form as the NtoN construction, we consider them to be near minimal pairs, following Weissweiler et al. (2022) who extract near minimal pairs from corpus data based on part-of-speech patterns. While these sentences inevitably contain more lexical biases than a true minimal pair dataset, they are completely natural, and provide a good comparison point for a construction where creating true minimal pairs is otherwise difficult (similar to the struggles of Weissweiler et al. (2022) regarding the Comparative Correlative construction). In total, we collect 456 total instances of NtoN distractors from COCA.
@@ 4.1 Methodology
We provide an additional, non-contextual baseline by training a linear classifier on GloVe embeddings for the nouns in the construction as input. It is well known that the NPN construction is biased towards certain lexical types of nouns, such as temporal phrases and body parts (Jackendoff, 2008). Thus, we expect that a classifier trained on the static embedding of the noun alone will achieve nontrivial performance. We argue that if a BERT-based classifier substantially outperforms this baseline, the difference is in performance is an indication of nontrivial contextual understanding of the construction as a whole, beyond the lexical semantics of the present nouns.
We train a separate probe based on embeddings from each layer of BERT and track performance across layers. We use the BERT-base-cased model, available through the Huggingface transformers library (Wolf et al., 2020), and choose logistic regression as our linear classification architecture.222We take the embedding of “to” as the input into the classifier, as some past work has considered it the “head” of the overall construction (Jackendoff, 2008). For all experiments and data settings, we run probes with 5 random seeds and report the average results.
@@ 4.2 Results
For the probing classifier results, we graph accuracy on the NtoN construction in Figure 1. As we can see, the classifier is relatively strong at distinguishing the NtoN construction from distractors even in the early layers, with an accuracy over .90 by layer 5 with full training examples. Additionally, the classifiers are robust to sharp reductions in the number of training examples (shown in lighter shades of green in Figure 1), showing strong performance even with as few as 10 per-class training examples. The control classifier achieves roughly chance performance, meaning that our trained probes have high selectivity (Hewitt and Liang, 2019). The lexical semantic baseline using GloVe achieves performance well above chance (68%), though it’s performance lags far behind the BERT-based probes, regardless of how many training example those BERT-based probes receive. This shows that overall, the probing classifier seems to be picking up on some sort of information in BERT which can reliably distinguish the NtoN construction from its near minimal pair NtoN distractor counterparts, beyond what is possible through lexical semantic clues alone. However, the distractor examples generally have syntactic structure which is divergent from the construction examples. To provide another comparison point, we now test if the existing probes can distinguish true instances of the NtoN construction from examples with artificially altered word orders.
@@ 5 Experiment 2: Perturbing Word Order
As we have seen in section 4.2, a BERT-based probe can generally distinguish the NtoN distractor patterns from the NtoN construction. However, we wish to further test how robust the model is at distinguishing the construction from related patterns. While we have compared to naturally occuring near minimal pairs, we now test the classifier on a set of examples artificially perturbed word order. If the classifier is robust at recognizing the NtoN construction, it should be able to correctly distinguish construction instances from artificial sentences with altered non-NPN word orders. To illustrate this point, consider the following two sentences:
@@ 5.1 Results
Figure 2 shows the probe’s performance on the perturbed test sets for the NtoN construction. Looking at Figure 2, we see that in the very early layers (1-3), the probe often predicts the NtoN construction despite the word order shifts, leading to relatively low accuracy. This possibly means that the classifier is biased by the lexical cues in the sentence early on, though performance is unstable between layers. Accuracy tends to peak in the later layers, with reduction in training examples leading to substantial drops in performance.
We find that there is some variation in performance depending on the specific artificial word order employed. Interestingly, performance on PN and PN perturbations is substantially worse in the earlier layers, though they eventually match or overtake the NP and NNP performance by the later layers. For NP and NNP, the models tend to learn the distinction quicker (by layer 4), while performance on PN and PNN does not stabilize to layers 7-8. By the late layers, most setting achieve respectable performance, besides NP/NNP which lag behind substantially when only trained on 10 examples.
@@ 5.2 Analysis
Overall, we find that classifier probes are able to distinguish instances of the NtoN construction from both near minimal pairs (NtoN distractor patterns) and artificial examples (perturbed word orderings). This work provides further evidence for the abilities of LMs to recognize the formal properties of a variety of constructions, as shown in previous work on other constructions (Li et al., 2022; Weissweiler et al., 2022; Mahowald, 2023). The peak in performance in the late-middle layers is consistent with much previous work on linguistic probing, which show that the middle and late-middle layers perform best for a variety of linguistic tasks (Goldberg, 2019; Hewitt and Manning, 2019; Lin et al., 2019; Liu et al., 2019).
The differences in the performance between the NP/NNP and the PN/PNN perturbed orderings is an unexpected finding. According to Rogers et al. (2021), the earlier layers of BERT encode “word order”, while the middle layers are where synactic capabilities emerge. Based on this logic, it is unsurprising that the classifier’s ability to distinguish PN/PNN emerges in the middle and later layers. Why might the NP/NNP instances be distinguished so much quicker? Our intuition is that in general, preposition tokens probably attend more to their immediately following word than their immediately preceding word. This is because prepositions are often immediately followed by objects, while their syntactic governor may or may not be directly adjacent to them. Perhaps in the early layers of the model (before hierarchy is as explicitly represented) prepositions learn to attend to their following token more quickly because this is a surface word order pattern that feeds quite well into syntax.
@@ 6.3 Results
Across all layers, both succession and juxtaposition perform worse with only 10 training examples, but performance stabilizes after only 25 examples for the probe. The relatively low recall for juxtaposition and succession when the classifiers are only trained with 10 examples indicates that the probe has not fully learned to correctly distinguish the two main semantic subtypes. It is somewhat striking that there is not a larger difference between succession and juxtaposition in performance, given that succession accounts for roughly 68% of all instances of the construction in our dataset. While probes are trained with balanced training sets, the relative frequency of these semantic subtypes within our dataset (and by extension COCA) is a strong indication that succession is the more frequent meaning. Nevertheless, performance is roughly comparable between the two semantic subtypes. In all cases, the distractor class is overpredicted, leading to a relatively low precision compared to the subtypes of the construction. As expected, the control classifiers achieve roughly chance performance across layers, indicating that our probes have high selectivity. The GloVe-based baseline achieves an average recall of around .54 across the subtypes, but has widely variable performance depending on the semantic subtype. In general, the GloVe based classifier is much more likely to underpredict succession, leading to very high precision and very low recall for this class.333We report GloVe and control results using the full training set. Performance of the GloVe baselines degrades with fewer examples, while the control classifiers remain near chance.
@@ 7 Related Work
Earlier computational linguistic work on English trained classifiers for such grammatico-semantic phenomena as identifying argument structure constructions (Hwang and Palmer, 2015) and disambiguating functions of tense and definiteness (Reichart and Rappoport, 2010; Bhatia et al., 2014), as well as generally to disambiguate the senses of prepositions (Litkowski and Hargraves, 2007; Schneider et al., 2018). Tayyar Madabushi et al. (2020) were the first to investigate BERT’s performance on learning constructions, finding that BERT is able to identify a large set of hundreds of automatically identified constructions. Regarding well-established argument structure constructions, Li et al. (2022) find that RoBERTa implicitly contains abstract knowledge of the constructions beyond specific lexical cues. Weissweiler et al. (2022) find that BERT-scale models are able to correctly distinguish the comparative-correlative construction from similar looking patterns, but find that the models fail on reasoning tests related to the construction’s semantics. Mahowald (2023) finds that the larger GPT-3 model can provide acceptability judgments for the Article+Adjective+Numeral+Noun (AANN) construction which generally align with human judgements, and find that the model is sensitive to constraints on the slots in the construction. Chronis et al. (2023) test BERT’s knowledge of the same AANN construction by projecting tokens in the construction into an interpretable embedding space, finding that features aligning with measure-words are evoked by tokens in the construction. Beyond BERT-scale models, Zhou et al. (2024), Bonial and Tayyar Madabushi (2024) and Scivetti et al. (2025) all test LLM knowledge of constructions in more complex scenarios, finding that their performance generally lags behind humans regarding construction understanding, though there is variation depending on the construction. Zhou et al. (2024) test a range of LLMs on understanding the causal-excess constructions in comparison to constructions with highly similar forms, showing that the model is often misled by form-based cues. Their experiments most closely mirror our inquiries into construction sense disambiguation, though they disambiguate between similar but distinct constructions while we focus on a single polysemous construction.

### v2

@@ Abstract
Construction Grammar hypothesizes that knowledge of a language consists chiefly of knowledge of form–meaning pairs (“constructions”) that include vocabulary, general grammar rules, and even idiosyncratic patterns. Recent work has shown that transformer language models represent at least some constructional patterns, including ones where the construction is rare overall. In this work, we probe BERT’s representation of the form and meaning of a minor construction of English, the NPN (noun–preposition–noun) construction—exhibited in such expressions as face to face and day to day—which is known to be polysemous. We construct a benchmark dataset of semantically annotated corpus instances (including distractors that superficially resemble the construction). With this dataset, we train and evaluate probing classifiers. They achieve decent discrimination of the construction from distractors, as well as sense disambiguation among true instances of the construction, revealing that BERT embeddings carry indications of the construction’s semantics. Moreover, artificially permuting the word order of true construction instances causes them to be rejected, indicating sensitivity to matters of form. We conclude that BERT does latently encode at least some knowledge of the NPN construction going beyond a surface syntactic pattern and lexical cues.111Code and Data: https://github.com/WesScivetti/NPN_probing
@@ 1 Introduction
Thus, there is need to evaluate language models on a range of linguistic tasks which go beyond the more studied “core” linguistic phenomena. Such work serves to provide a more complete picture of how language models succeed and fail across the broad spectrum of phenomena in language. Indeed, beyond the more mainstream notions of linguistic structure and information, there is also work on investigating LM knowledge of more idiosyncratic constructions, as defined by Construction Grammar. Construction Grammar is broadly a family of linguistic theories which consider all parts of language to be made up of constructions, which are pairings of linguistic forms with meaning or function (Goldberg 1995; Croft 2001, inter alia). It remains unclear the extent to which LMs may implicitly view constructions as distinct units. Because of their emphasis on pairing form with meaning, CxG theories provide possibilities for testing language model capabilities at the interface of form and meaning for different aspects of language, in contrast to past work which has focused on either syntax (e.g. Hewitt and Manning 2019) or semantics (e.g. Vulić et al. 2020) in isolation. A substantial and growing amount of research has recently focused on the intersection of LM knowledge and Construction Grammar (Tayyar Madabushi et al., 2020; Tseng et al., 2022; Pannitto and Herbelot, 2023; Veenboer and Bloem, 2023, inter alia), with a particular focus on argument structure constructions (Li et al., 2022), the English Comparative Correlative (Weissweiler et al., 2022), and the English AANN construction (Chronis et al., 2023; Mahowald, 2023). While these studies have provided valuable insight into LM processing of constructions with varying levels of schematicity, there remain many constructions which have not been addressed at all in previous work. Furthermore, while Zhou et al. (2024) do test model understanding of constructions which are similar in form, no past work has focused on individual constructions as polysemous units. We argue this is a gap in past work, as constructions, like words, can have related but distinct meanings that must be properly disambiguated in context in order for correct interpretation. We address this gap by providing experiments which pair formal sensitivity with semantic disambiguation in a controlled manner for a single construction.
@@ 2 The NPN Construction
Given the general rules of English, the NPN construction has several unique properties, which we argue separate it from more “core” linguistic phenomena. Firstly, the nouns almost always lack determiners, which is unusual for count nouns like “day”. Secondly, the construction can occur in a variety of syntactic positions, including as an adverbial modifier (as in section 2) and as a prenominal modifier (as in section 2). Finally, it conveys a meaning which is not entirely predictable from its components, and varies considerably depending on the preposition. Common meanings of the NPN construction are the succession meaning (shown in section 2) and the matching/comparison meaning (shown in section 2). See Jackendoff (2008) for an overview of the NPN construction and the common meanings associated with various prepositional lemmas.
@@
While there are additional meanings of NPN that do not occur with “to” as the preposition, it is one of the only prepositions that is ambiguous in the NPN construction. By not considering examples of NPN with other prepositions, we remove the prepositional lemma as a potential shallow cue that models could learn to predict the construction’s semantics. While there are arguably examples of NPNs where the two nouns are not identical, we limit our analysis to cases where the two nouns in the construction match exactly. This allows us to easily gather examples of the construction from corpus data.
@@ 3.1 Corpus Gathering and Cleaning
After extracting all sentences which contained a possible instance of NtoN, we then manually clean the data, removing sentences that were either too short (<5 tokens) or contained too many typos. We annotate all instances of the construction for their semantic subtype, and double annotate roughly 25% of the dataset, achieving an agreement of 84% and a Cohen’s kappa value of .754 between the two annotators, indicating strong agreement. 222Disagreements between the two annotators were resolved through discussion and a gold label was chosen jointly. The final dataset has 6599 instances of NtoN, of which 1885 were double annotated.
@@ 3.2 Near Minimal Pairs
We do not consider such cases to be examples of the NtoN construction because the surface pattern of Noun + Preposition + Noun clearly arises from a different syntactic context (e.g. a verb licensing a direct object and a PP modifier). Furthermore, the meanings of these examples do not evoke the unique semantics that accompany the NtoN construction. While these cases are not instances of the NtoN construction, they do provide a set of negative examples which we can use to probe the model’s ability to recognize true NtoN constructions. Throughout this paper, we refer to this set of examples as instances of the NtoN distractors, since we test of if the model is “distracted” by the shallow similarity of the examples to the NPN construction. We refer to true examples of NtoN as instances of the NtoN construction. Since these NtoN examples exhibit the same surface form as the NtoN construction, we consider them to be near minimal pairs, following Weissweiler et al. (2022) who extract near minimal pairs from corpus data based on part-of-speech patterns. While these sentences inevitably contain more lexical biases than a true minimal pair dataset, they are completely natural, and provide a good comparison point for a construction where creating true minimal pairs is otherwise difficult (because there is no obvious minimal change that can be made to result in a grammatical sentence that is not an example of the construction, similar to the struggles of Weissweiler et al. (2022) regarding the Comparative Correlative construction). In total, we collect 456 total instances of NtoN distractors from COCA.
@@ 4.1 Methodology
We provide an additional, non-contextual baseline by training a linear classifier on GloVe embeddings for the nouns in the construction as input. It is well known that the NPN construction is biased towards certain lexical types of nouns, such as temporal phrases and body parts (Jackendoff, 2008). Thus, we expect that a classifier trained on the static embedding of the noun alone will achieve nontrivial performance. We argue that if a BERT-based classifier substantially outperforms this baseline, the difference in performance is an indication of nontrivial contextual understanding of the construction as a whole, beyond the lexical semantics of the present nouns.
Following previous probing work which tracks performance layer by layer Liu et al. (2019); Weissweiler et al. (2022), we train a separate probe based on embeddings from each layer of BERT and track performance across layers. We use the BERT-base-cased model, available through the Huggingface transformers library (Wolf et al., 2020), and choose logistic regression as our linear classification architecture.333We take the embedding of “to” as the input into the classifier, as some past work has considered it the “head” of the overall construction (Jackendoff, 2008). For all experiments and data settings, we run probes with 5 random seeds and report the average results.
@@ 4.2 Results
For the probing classifier results, we graph accuracy on the NtoN construction in Figure 1. As we can see, the classifier is relatively strong at distinguishing the NtoN construction from distractors even in the early layers, with an accuracy over .90 by layer 5 with full training examples. Additionally, the classifiers are robust to sharp reductions in the number of training examples (shown in lighter shades of green in Figure 1), showing strong performance even with as few as 10 per-class training examples, echoing similar findings for other constructions (Tayyar Madabushi et al., 2020). The control classifier achieves roughly chance performance, meaning that our trained probes have high selectivity (Hewitt and Liang, 2019). The lexical semantic baseline using GloVe achieves performance well above chance (68%), though its performance lags far behind the BERT-based probes, regardless of how many training example those BERT-based probes receive. This shows that overall, the probing classifier seems to be picking up on some sort of information in BERT which can reliably distinguish the NtoN construction from its near minimal pair NtoN distractor counterparts, beyond what is possible through lexical semantic clues alone. However, the distractor examples generally have syntactic structure which is divergent from the construction examples. To provide another comparison point, we now test if the existing probes can distinguish true instances of the NtoN construction from examples with artificially altered word orders.
@@ 5 Experiment 2: Perturbing Word Order
As we have seen in section 4.2, a BERT-based probe can generally distinguish the NtoN distractor patterns from the NtoN construction. However, we wish to further test how robust the model is at distinguishing the construction from related patterns. While we have compared to naturally occuring near minimal pairs, we now test the classifier on a set of examples with artificially perturbed word order. If the classifier is robust at recognizing the NtoN construction, it should be able to correctly distinguish construction instances from artificial sentences with altered non-NPN word orders. To illustrate this point, consider the following two sentences:
@@ 5.1 Results
Figure 2 shows the probe’s performance on the perturbed test sets for the NtoN construction. We see that in the very early layers (1–3), the probe often predicts the NtoN construction despite the word order shifts, leading to relatively low accuracy. This possibly means that the classifier is biased by the lexical cues in the sentence early on. Interestingly, performance on PN and PN perturbations is substantially worse than performance on NP and NNP in the early layers. Accuracy on all perturbations trends upwards in the later layers, with reduction in training examples leading to drops in performance especially for NP/NNP.
@@ 5.2 Analysis
Overall, we find that classifier probes are able to distinguish instances of the NtoN construction from both near minimal pairs (NtoN distractor patterns) and artificial examples (perturbed word orderings). This finding aligns with the strong performance on form-based recognition that has been observed in previous work on other constructions (Li et al., 2022; Weissweiler et al., 2022; Mahowald, 2023). The peak in performance in the late-middle layers is consistent with much previous work on linguistic probing, which show that the middle and late-middle layers perform best for a variety of linguistic tasks (Goldberg, 2019; Hewitt and Manning, 2019; Lin et al., 2019; Liu et al., 2019).
The differences in the performance between the NP/NNP and the PN/PNN perturbed orderings is an unexpected finding. According to Rogers et al. (2021), the earlier layers of BERT encode “word order”, while the middle layers are where syntactic capabilities emerge. Based on this logic, it is unsurprising that the classifier’s ability to distinguish PN/PNN emerges in the middle and later layers. Why might the NP/NNP instances be distinguished so much quicker? Our intuition is that in general, preposition tokens probably attend more to their immediately following word than their immediately preceding word. This is because prepositions are often immediately followed by objects, while their syntactic governor may or may not be directly adjacent to them. Perhaps in the early layers of the model (before hierarchy is as explicitly represented) prepositions attend to their following token more quickly because this is a surface word order pattern that feeds quite well into syntax.
@@ 6.3 Results
Across all layers, both succession and juxtaposition perform worse with only 10 training examples, but performance stabilizes after only 25 examples for the probe. The relatively low recall for juxtaposition and succession when the classifiers are only trained with 10 examples indicates that the probe has not fully learned to correctly distinguish the two main semantic subtypes. It is somewhat striking that there is not a larger difference between succession and juxtaposition in performance, given that succession accounts for roughly 68% of all instances of the construction in our dataset. While probes are trained with balanced training sets, the relative frequency of these semantic subtypes within our dataset (and by extension COCA) is a strong indication that succession is the more frequent meaning. Nevertheless, performance is roughly comparable between the two semantic subtypes. In all cases, the distractor class is overpredicted, leading to a relatively low precision compared to the subtypes of the construction. As expected, the control classifiers achieve roughly chance performance across layers, indicating that our probes have high selectivity. The GloVe-based baseline achieves an average recall of around .54 across the subtypes, but has widely variable performance depending on the semantic subtype. In general, the GloVe based classifier is much more likely to underpredict succession, leading to very high precision and very low recall for this class.444We report GloVe and control results using the full training set. Performance of the GloVe baselines degrades with fewer examples, while the control classifiers remain near chance.
@@ 7 Related Work
Earlier computational linguistic work on English trained classifiers for such grammatico-semantic phenomena as identifying argument structure constructions (Hwang and Palmer, 2015) and disambiguating functions of tense and definiteness (Reichart and Rappoport, 2010; Bhatia et al., 2014), as well as generally to disambiguate the senses of prepositions (Litkowski and Hargraves, 2007; Schneider et al., 2018). Tayyar Madabushi et al. (2020) were the first to investigate BERT’s performance on learning constructions, finding that BERT is able to identify a large set of hundreds of automatically identified constructions. Regarding well-established argument structure constructions, Li et al. (2022) find that RoBERTa implicitly contains abstract knowledge of the constructions beyond specific lexical cues. Weissweiler et al. (2022) find that BERT-scale models are able to correctly distinguish the comparative-correlative construction from similar looking patterns, but find that the models fail on reasoning tests related to the construction’s semantics. Mahowald (2023) finds that the larger GPT-3 model can provide acceptability judgments for the Article+Adjective+Numeral+Noun (AANN) construction which generally align with human judgements, and find that the model is sensitive to constraints on the slots in the construction. Chronis et al. (2023) test BERT’s knowledge of the same AANN construction by projecting tokens in the construction into an interpretable embedding space, finding that features aligning with measure-words are evoked by tokens in the construction. Beyond BERT-scale models, Zhou et al. (2024), Bonial and Tayyar Madabushi (2024) and Scivetti et al. (2025) all test LLM knowledge of constructions in more complex scenarios, finding that their performance generally lags behind humans regarding construction understanding, though there is variation depending on the construction. Zhou et al. (2024) test a range of LLMs on understanding the causal-excess constructions in comparison to constructions with highly similar forms, showing that the model is often misled by form-based cues. Their experiments most closely mirror our inquiries into construction sense disambiguation, though they disambiguate between similar but distinct constructions while we focus on a single polysemous construction. While Zhou et al. (2024) find that LLMs largely are unsucessful at meaning-based disambiguation, and Weissweiler et al. (2022) also find negative results regarding the semantics of the comparative-correlative, our relatively positive results on construction disambiguation in this present work demonstrate that for NtoN, models may possess more robust models of constructional semantics than would be previously expected.
While NPN has not been the major focus of past analysis, Weissweiler et al. (2024) do consider it as one of the constructions which they include in their UCxn dataset, which is compiled by automatically using Universal Dependencies (de Marneffe et al., 2021) graphs to find indications of constructions across 10 languages. We do not use this dataset due to its limited size (it contains under 50 total examples of the NPN construction in English).